OpenStudy (anonymous):

For questions 4-5, how many solutions does the system have? 4. 4y+4=x 2x+8y=8 a. one solution b. two solutions c. infinitely many solutions d. no solution 5. 6x+2=y 3y-18x=12 a. one solution b. two solutions c. infinitely many solutions d. no solution

OpenStudy (hoblos):

4y+4=x 2x+8y=8 if we divide the second equation by we'll get x+4y = 4 then x=4-4y replace it in the first equation 4y+4 = 4-4y 8y=0 then y=0 x=4 thus this system has one solution

OpenStudy (anonymous):

so #4 has one solution?

OpenStudy (hoblos):

6x+2=y 3y-18x=12 replace y from first equation in the second 3(6x+2) - 18x = 12 18x +6 -18x = 12 6=12 impossible! then the system has no solutions

OpenStudy (hoblos):

yeah #4 has one solution #5 has no solution

OpenStudy (anonymous):

ok thank you so much i couldnt understand any of this

OpenStudy (hoblos):

glad to help :)

OpenStudy (anonymous):

the answer to #4 was actually many soltuions but thats ok.

OpenStudy (anonymous):

System 4: One intersection of the 2 lines, meaning 1 solution. System 5: The 2 lines are parallel (have the same slope). This means an infinity of solutions.